Vitex, Inc. manufactures a popular consumer product and it has provided the following data excerpts from its standard cost system:

Inputs (1) Standard Quantity or Hours (2)
Standard
Price
or Rate
Standard
Cost
(1) × (2)
Direct materials 2.20 pounds $ 16.40 per pound $ 36.08
Direct labor 1.00 hours $ 15.70 per hour $ 15.70
Variable manufacturing overhead 1.00 hours $ 9.60 per hour $ 9.60
Total standard cost per unit $ 61.38
Total Variances Reported
Standard
Cost*
Price
or Rate
Quantity or
Efficiency
Direct materials $ 577,280 $ 10,788 F $ 32,800 U
Direct labor $ 251,200 $ 3,400 U $ 15,700 U
Variable manufacturing overhead $ 153,600 $ 4,700 F $ ? U

*Applied to Work in Process during the period.

The company's manufacturing overhead cost is applied to production on the basis of direct labor-hours. All of the materials purchased during the period were used in production. Work in process inventories are insignificant and can be ignored.

Required:

1. How many units were produced last period?

2. How many pounds of direct material were purchased and used in production?

3. What was the actual cost per pound of material? (Round your answer to 2 decimal places.)

4. How many actual direct labor-hours were worked during the period?

5. What was the actual rate paid per direct labor-hour? (Round your answer to 2 decimal places.)

6. How much actual variable manufacturing overhead cost was incurred during the period?
Answer #1
1. Computation of Actual Unit Produced
Standard material cost (a) 577280
Standard material cost per unit (b) $36
Actual Unit Prodcued ( a/b)                                  16,000.00
2. Computation of Pound Purchased and used
Std Material quantity per unit (a) 2.20 Pound
Actual Unit prodcued (b) 16000
Std Quantity allowed for actual output (c=Axb) 35200
Material Quantity Variance=$32800 U
-32800= ( SQ-AQ)X STd Price
-32800=(35200-AQ)X 2.20
AQ=50110 Poound
Computation of Actual Cost per Pound
Material Price variance=10788 F
10788 = Actual Quantity (Std price -Actual price)
10788 = 50110 X (2.20- Actual Price)
Actual price per pound = $ 1.9847 per pound
Computation of Actual Labour Hour Worked
Std Labour hours for actual output = 16000*1 = 16,000 hours
Std labour rate per hour = $ 15.70 per hour
Labour efficiency variance= 15700 U
Labour efficiency variancce= Std rate per hour (Std hours -Actual hours)
-15700 = 15.70 (16000- Actual hours worked)
Actual Hours worked = 17000 hours
Computation of Actual Rate per Hour
Labour rate variance= 3400 U
Labour rate variance= Actual hours (Std rate -Actual rate)
-3400 = 17000 hours (15.70-Actual rate er hour)
Actual rate per hour: $ 15.90 per hour
Computation of Actual Variable Manufacturing OH incurred
Std labour hours: 16000 hours
Actual hours: 17000 hours
Std rate per hour: $ 9.60 per hour
Variable OH rate variance = (Std OH rate*Actual hours - Actual variable OH incurred)
4700 F = (17000 hours*9.60) - Actual variable OH incurred
Actual variable OH incurred= $158500